OpenStudy (anonymous):

Identify the degree of the polynomial 4c^3d^2 + 7cd^3 – 11c^2d + 1

OpenStudy (anonymous):

is it 3?

sam (.sam.):

Nope

sam (.sam.):

For the exponent in 4c^3d^2, you'll need to add them up

OpenStudy (mathstudent55):

Take each term and look at its variables. Add the exponents of all the variables of each term. The degree of each term is the sum of the exponents of the variables of the term. The degree of the polynomial is the same as the highest degree any term.

OpenStudy (anonymous):

5?

OpenStudy (mathstudent55):

Yes

OpenStudy (mathstudent55):

4c^3d^2 + 7cd^3 – 11c^2d + 1 ^ ^ ^ ^ ^ ^ ^ 3+2 = 5 1 + 3 = 4 2 + 1 = 3 0 The degrees of the ternms are 5, 4, 3, and 0 The highest degree of any term is 5, so the degree of the polynomial is 5. Keep in mind that a variable with no exponent has an exponent of 1, and a number by itself with no variable has degree 0.

OpenStudy (anonymous):

Identify the degree of the polynomial 5h4 – 22h2 + 9

OpenStudy (mathstudent55):

By now you should be able to do this.

OpenStudy (mathstudent55):

What is the degree of each term?

OpenStudy (anonymous):

im confused